# Ask in Any Modality: A Comprehensive Survey on Multimodal Retrieval-Augmented Generation

### v1

@@ 3 Evaluation
Assessing multimodal RAG models presents significant challenges due to their diverse modalities, intricate architectures, and the diverse evaluation dimensions involved. Given the broad spectrum of modalities and system components, evaluation requires a combination of metrics drawn from vision-language models (VLMs), generative AI, and retrieval systems. These metrics enable the assessment of various aspects of multimodal RAG, such as text and image generation or information retrieval, either independently or in combination. By analyzing prior works, we identified approximately 60 distinct evaluation metrics used in multimodal RAG studies. More details, including the formulas for the RAG evaluation metrics, can be found in Appendix C. In the following paragraphs, we will examine the most important and widely used metrics for evaluating multimodal RAG.
@@
Evaluates how well a system retrieves relevant information from a database or knowledge source. Accuracy, recall, and precision are used to evaluate retrieval and downstream task performance, the F1 score also being used to assess both Recall and Precision simultaneously. A more commonly used metric than recall is Recall@K, which measures the proportion of relevant items appearing within the top K retrieved results. Another important metric Mean Reciprocal Rank (MRR), which is utilized by Adjali et al. (2024); Nguyen et al. (2024).
@@
These evaluations focus on criteria that rely on modalities. Text and image are the most commonly used modalities in multimodal RAG. This category of evaluations considers various aspects of these modalities, including the alignment between generated text and images, the fluency of the text, the quality of image captions, and other related factors. In text similarity and overlap metrics, Exact Match (EM), BLEU Papineni et al. (2002), ROUGE Lin (2004) and METEOR Banerjee and Lavie (2005) are used. MultiRAGen Shohan et al. (2024) uses Multilingual ROUGE for multilingual settings. For image captioning evaluation, CIDEr (Consensus-Based Image Description Evaluation) Vedantam et al. (2015) is a widely used metric for evaluating image captioning quality. It quantifies the alignment between a generated caption and a set of human-written reference captions by leveraging the term frequency-inverse document frequency (TF-IDF) and cosine similarity. Yasunaga et al. (2023); Zhao et al. (2024); Luo et al. (2024a); Yuan et al. (2024); Sharifymoghaddam et al. (2024); Hu et al. (2023); Rao et al. (2024); Xu et al. (2024a); Kim et al. (2024); Zhang et al. (2024c) use CIDEr for evaluation. SPICE (Semantic Propositional Image Caption Evaluation) Anderson et al. (2016) is a metric that emphasizes semantic aspects, in contrast to CIDEr. SPIDEr Liu et al. (2017), utilized in Zhang et al. (2024c), is a linear combination of both SPICE and CIDEr. In semantic alignment evaluations, BERTScore Zhang et al. (2020), employed in Sun et al. (2024b); Shohan et al. (2024), measures semantic similarity by comparing BERT-generated embeddings. Chen et al. (2022a); Zhi Lim et al. (2024); Ma et al. (2024c) evaluate Fluency by employing BERTScore. CLIP Score Hessel et al. (2021), used in Sharifymoghaddam et al. (2024); Zhang et al. (2024c), measures image-text similarity using the CLIP model Radford et al. (2021). FID (Fréchet Inception Distance) Heusel et al. (2017) quantifies image quality by computing the distributional difference between the generated and original image features. Yasunaga et al. (2023); Zhao et al. (2024); Sharifymoghaddam et al. (2024); Zhang et al. (2024c) use FID in their works. Zhao et al. (2024) use Kernel Inception Distance (KID) Bińkowski et al. (2018) for evaluation. KID is another metric used to evaluate the quality of generated images. It is similar to FID but offers an unbiased estimator. Zhi Lim et al. (2024) employs the Inception Score (IS) to assess the diversity and quality of generated images based on classification probabilities obtained from the Inception network. For the evaluation of audio synthesis, Zhang et al. (2024c) assesses overall sound quality (OVL) and relevance to the text (REL) with the assistance of human annotators. Additionally, this study employs the Fréchet Audio Distance (FAD) Kilgour et al. (2019) as an evaluation metric. FAD serves a similar purpose to the FID but is specifically designed for the audio domain.
FLOPs, execution time, response time, and average retrieval time per query are utilized to assess efficiency and computational performance Nguyen et al. (2024); Strand et al. (2024); Dang (2024); Zhou (2024). Zhong et al. (2024) employs Spearman’s Rank Correlation (SRC), which calculates the Pearson correlation between the rank values of two variables. Several evaluation metrics are domain- and problem-specific. Geodesic distance measurement quantifies the model’s accuracy by calculating the geodesic distance between predicted and actual geographical coordinates Zhou et al. (2024e). Lahiri and Hu (2024) utilizes Clinical Relevance (CR) for evaluation.
@@ 4.1 Retrieval Strategy
Efficient Search and Similarity Retrieval
@@
Modern multimodal RAG systems first encode diverse modalities (text, images, tables) into a unified embedding space to enable direct cross-modal retrieval. The foundational CLIP-Radford et al. (2021) or BLIP-based Li et al. (2022a) approaches leverage contrastive learning, which aligns text and images into a shared latent space via dual-tower encoders. Recent advancements extend this paradigm with novel multimodal retrieval architectures and training strategies Zhou et al. (2024c); Wei et al. (2025); Zhang et al. (2024j). Multimodal RAG systems rely on efficient search strategies to retrieve relevant external knowledge from large-scale multimodal memory. Since dual-encoder models map multimodal inputs into a shared high-dimensional embedding space, maximum inner product search (MIPS) is commonly used for fast and direct similarity comparisons. Systems such as MuRAG Chen et al. (2022a) and RA-CM3 Yasunaga et al. (2023) employ approximate MIPS to efficiently retrieve the top-k candidates by maximizing the inner product between the query vector and a large collection of image–text embeddings. Large-scale implementations leverage distributed MIPS techniques, such as TPU-KNN Chern et al. (2022), to enable high-speed retrieval across extensive datasets. Additionally, ScaNN (Scalable Nearest Neighbors) Guo et al. (2020), MAXSIM score Chan and Ng (2008); Cho et al. (2024), and approximate KNN methods Caffagni et al. (2024) have been adopted for efficient image-text similarity computation.
@@
Recent advancements in MIPS optimization focus on reducing retrieval latency and improving accuracy. These include adaptive quantization Li et al. (2024a); Huang et al. (2024b), hybrid sparse-dense representations Nguyen et al. (2024); Ma et al. (2024a); Zhang et al. (2024a), and learned index structures Zhai et al. (2023); Basnet et al. (2024). These techniques help optimize retrieval performance by balancing efficiency and precision in multimodal RAG systems.
Modality-Based Retrieval
@@
Modality-aware retrieval techniques optimize retrieval efficiency by leveraging the unique characteristics of each modality. Text-centric retrieval remains foundational in multimodal RAG systems. Traditional methods like BM25 Robertson and Zaragoza (2009), along with dense retrievers such as BGE-1.5 Xiao et al. (2023) and BGE-M3 Chen et al. (2024a) dominate text-based evidence retrieval Chen et al. (2022b); Suri et al. (2024); Nan et al. (2024). Novel approaches also address the need for fine-grained semantic matching and domain specificity: for instance, ColBERT Khattab and Zaharia (2020) and PreFLMR Lin et al. (2024b) employ token-level interaction mechanisms that preserve nuanced textual details to improve precision for multimodal queries, while RAFT Zhang et al. (2024i) and CRAG Yan et al. (2024) enhance retrieval performance by ensuring precise citation of text spans, thereby optimizing domain-specific question answering within multimodal RAG pipelines. Vision-centric retrieval focuses on directly leveraging image representations for knowledge extraction Kumar and Marttinen (2024); Yuan et al. (2023). Systems such as EchoSight Yan and Xie (2024) and ImgRet Shohan et al. (2024) retrieve visually similar content by using reference images as queries. Composed Image Retrieval (CMI) models Feng et al. (2023); Zhao et al. (2024); Jang et al. (2024) enhance retrieval by integrating multiple image features into a unified query representation. Similarly, Pic2word Saito et al. (2023) maps visual content to textual descriptions, enabling zero-shot image retrieval. Video-centric retrieval extends vision-based techniques by incorporating temporal dynamics and large video-language models (LVLMs), driven by novel frameworks like iRAG Arefeen et al. (2024), which introduces incremental retrieval for sequential video understanding, and MV-Adapter Jin et al. (2024), optimizing multimodal transfer learning for video-text retrieval. Recent breakthroughs focus on long-context processing: Video-RAG Luo et al. (2024b) leverages visually aligned auxiliary texts (OCR/ASR) to enhance retrieval without proprietary models, while VideoRAG Ren et al. (2025) employs dual-channel architectures with graph-based knowledge grounding for extreme-length videos. For temporal reasoning, CTCH Shen et al. (2024) uses contrastive transformer hashing to model long-term dependencies, and RTime Du et al. (2024) introduces reversed-video hard negatives to rigorously benchmark temporal causality. Meanwhile, for complex video understanding, OmAgent Zhang et al. (2024e) adopts a divide-and-conquer framework, while DRVideo Ma et al. (2024d) addresses long-video understanding by a document-based retrieval approach.
@@
Recent research has advanced beyond traditional uni-modal retrieval methods toward models that directly process entire document pages. These approaches integrate textual, visual, and structural elements like tables, images, font styles, and page layouts to enhance retrieval performance in complex documents. ColPali Faysse et al. (2024) pioneers end-to-end document image retrieval by embedding page patches with a vision-language backbone, bypassing OCR entirely. Models like ColQwen2 Wang et al. (2024b); Khattab and Zaharia (2020) and M3DocVQA Cho et al. (2024) extend this paradigm with dynamic resolution handling and holistic multi-page reasoning. Newer frameworks refine efficiency and layout understanding: ViTLP Mao et al. (2024) and DocLLM Wang et al. (2024a) pre-train generative models to align spatial layouts with text, while CREAM Zhang et al. (2024b) employs coarse-to-fine retrieval with multimodal efficient tuning to balance accuracy and computation costs. Finally, mPLUG-DocOwl 1.5 Hu et al. (2024a) and 2 Hu et al. (2024b) unify structure learning across formats (e.g., invoices, forms) without OCR dependencies. Together, these approaches highlight a shift toward holistic, layout-sensitive retrieval that leverages visual semantics and generative pre-training.
Re-ranking and Selection Strategies
Effective retrieval in multimodal RAG systems requires not only identifying relevant information but also refining and prioritizing retrieved candidates. Re-ranking and selection strategies enhance retrieval quality by optimizing example selection, refining relevance scoring, and applying filtering mechanisms. Optimized example selection techniques often employ multi-step retrieval, integrating both supervised and unsupervised selection strategies Luo et al. (2024a). Su et al. (2024a) refine multimodal inputs using probabilistic control keywords to improve credibility, while RULE Xia et al. (2024b) calibrates retrieved context selection through statistical methods like the Bonferroni correction to mitigate factuality risks. Furthermore, selective fusion mechanisms dynamically refine the retrieval outcomes Yuan et al. (2023), and key-frame extraction techniques ensure diversity in video-based retrieval through clustering-based frame selection Dong et al. (2024b). Several methods employ advanced scoring mechanisms to improve retrieval relevance Mortaheb et al. (2025b, a). Multimodal similarity measures, including negative log-likelihood (NLL), structural similarity index measure (SSIM) Wang et al. (2020), normalized cross-correlation (NCC), and BERTScore Zhang et al. (2020), aid in re-ranking retrieved documents based on textual and visual similarity. Cross-encoders trained on sequence classification tasks refine relevance scoring within retrieval pipelines Zhi Lim et al. (2024), while hierarchical post-processing techniques integrate retrieval, passage-level, and answer confidence scores for improved ranking Zhang et al. (2024h). LDRE Yang et al. (2024) employs semantic ensemble methods to adaptively weigh multiple caption features, whereas RAGTrans Cheng et al. (2024) and OMG-QA Nan et al. (2024) integrate traditional ranking functions like BM25 Robertson and Zaragoza (2009). Advanced re-ranking techniques also incorporate Q-Former-based image-to-text correspondence for fine-grained retrieval Yan and Xie (2024); Xu et al. (2024a). Advanced filtering methods are crucial for enhancing multi-modal RAG performance by ensuring that only high-quality, relevant data is processed. Hard negative mining techniques, as seen in GME Zhang et al. (2024j) and MM-Embed Lin et al. (2024a), use modality-aware sampling and synthesized negatives from imbalanced data to mitigate modality bias and refine contrastive learning. Similarly, consensus-based filtering approaches, exemplified by MuRAR Zhu et al. (2024d) and ColPali Faysse et al. (2024), employ source attribution and multi-vector mapping to filter out low-similarity candidates and reduce hallucinations in enterprise QA. Moreover, dynamic modality filtering methods—including RAFT Zhang et al. (2024i), Img2Loc Zhou et al. (2024e), and MAIN-RAG Chang et al. (2024)—train retrievers to disregard confusing data, thereby enhancing the discriminative capacity and overall robustness of multi-modal retrieval systems.
@@ 4.2 Fusion Mechanisms
Score Fusion and Alignment
@@
Recent multimodal RAG models emphasize aligning embeddings or fusing scores across different modalities to enable adaptive retrieval. Zhi Lim et al. (2024) transform text, tables, and images into a unified textual representation using a cross-encoder trained for relevance scoring. To address multimodal LLM input constraints, Sharifymoghaddam et al. (2024) introduce interleaved image-text pairs, merging multiple few-shot images vertically for compatibility with models like LLaVA Liu et al. (2023a) while aligning modalities via CLIP score fusion Hessel et al. (2021) and BLIP feature fusionLi et al. (2022a). For fine-grained alignment, REVEAL Hu et al. (2023) injects retrieval scores into attention layers and minimizes L2-Norm distances between query and knowledge embeddings, and MA-LMM He et al. (2024) aligns video-text embeddings via a BLIP-inspired Query Transformer (Q-Former)Li et al. (2022a). In contrast, LLM-RA Jian et al. (2024) concatenates text and visual embeddings into joint queries to reduce retrieval noise. RA-BLIP Ding et al. (2024b) employs a 3-layer BERT-based adaptive fusion module to unify visual-textual semantics. CLIP-based alignment strategies are widely adopted; Riedler and Langer (2024) embed images and questions into a shared CLIP space, while VISA Ma et al. (2024b) uses the Document Screenshot Embedding (DSE) model as the retrieval component of the RAG system to align and fuse textual queries with visual document representations. DSE directly encodes both textual queries and document screenshots into a shared embedding space, allowing efficient retrieval based on cosine similarity. Xue et al. (2024) map object-predicate pairs into a shared semantic space using a prototype-based embedding network (PENET) Zheng et al. (2023), aligning visual features with textual prototypes. Re-IMAGEN Chen et al. (2022b) refines text-to-image synthesis with interleaved classifier-free guidance during diffusion sampling, balancing creativity and entity fidelity. To improve multimodal alignment, MegaPairs Zhou et al. (2024a) integrates CLIP-based and MLLM-based retrieval through score fusion. Meanwhile, Wiki-LLaVA Caffagni et al. (2024) employs CLIP encoders for hierarchical document retrieval. VISRAG Yu et al. (2024) refines alignment via position-weighted mean pooling on VLM hidden states, prioritizing later tokens for relevance. RAG-Driver Yuan et al. (2024) aligns visual-language embeddings via visual instruction tuning and an MLP projector.
Attention-Based Mechanisms

### v2

@@ 3 Evaluation
Evaluating multimodal RAG models is complex due to their varied input types and complex structure. The evaluation combines metrics from VLMs, generative AI, and retrieval systems to assess capabilities like text/image generation and information retrieval. Our review found about 60 different metrics used in the field. More details, including the formulas for the RAG evaluation metrics, can be found in Appendix (§C). In the following paragraphs, we will examine the most important and widely used metrics for evaluating multimodal RAG.
@@
Retrieval performance is measured through accuracy, recall, and precision metrics, with an F1 score combining recall and precision. Recall@K, which examines relevant items in top K results, is preferred over standard recall. Mean Reciprocal Rank (MRR) serves as another key metric for evaluation, which is utilized by Adjali et al. (2024); Nguyen et al. (2024).
@@
Modality-based evaluations primarily focus on text and image, assessing their alignment, text fluency, and image caption quality. For text evaluation, metrics include Exact Match (EM), BLEU Papineni et al. (2002), ROUGE Lin (2004), and METEOR Banerjee and Lavie (2005). MultiRAGen Shohan et al. (2024) uses Multilingual ROUGE for multilingual settings.
For image captioning, CIDEr (Consensus-Based Image Description Evaluation) Vedantam et al. (2015) measures caption quality using TF-IDF and cosine similarity Yasunaga et al. (2023); Zhao et al. (2024); Luo et al. (2024a); Yuan et al. (2024); Sharifymoghaddam et al. (2024); Hu et al. (2023); Rao et al. (2024); Xu et al. (2024a); Kim et al. (2024); Zhang et al. (2024c), while SPICE (Semantic Propositional Image Caption Evaluation) Anderson et al. (2016) focuses on semantics. SPIDEr Liu et al. (2017), used in Zhang et al. (2024c), combines both metrics. For semantic alignment, BERTScore Zhang et al. (2020) compares BERT embeddings Sun et al. (2024b); Shohan et al. (2024), and evaluates fluency Chen et al. (2022a); Zhi Lim et al. (2024); Ma et al. (2024c). CLIP Score Hessel et al. (2021), used in Sharifymoghaddam et al. (2024); Zhang et al. (2024c), measures image-text similarity using CLIP Radford et al. (2021). For image quality, FID (Fréchet Inception Distance) Heusel et al. (2017) compares feature distributions Yasunaga et al. (2023); Zhao et al. (2024); Sharifymoghaddam et al. (2024); Zhang et al. (2024c), while KID (Kernel Inception Distance) Bińkowski et al. (2018) provides an unbiased alternative. Inception Score (IS) evaluates image diversity and quality through classification probabilities Zhi Lim et al. (2024). For audio evaluation, Zhang et al. (2024c) uses human annotators to assess sound quality (OVL) and text relevance (REL), while also employing Fréchet Audio Distance (FAD) Kilgour et al. (2019), an audio-specific variant of FID.
System efficiency is measured through FLOPs, execution time, response time, and retrieval time per query Nguyen et al. (2024); Strand et al. (2024); Dang (2024); Zhou (2024). Domain-specific metrics include geodesic distance for geographical accuracy Zhou et al. (2024e), and Clinical Relevance for medical applications Lahiri and Hu (2024).
@@ 4.1 Retrieval Strategy
4.1.1 Efficient Search and Similarity Retrieval
Modern multimodal RAG systems encode diverse input modalities into a unified embedding space to enable direct cross-modal retrieval. Recent advancements in CLIP-based Radford et al. (2021) and BLIP-inspired Li et al. (2022a) approaches have driven the evolution of contrastive learning strategies through novel multimodal retrieval architectures and training methodologies Zhou et al. (2024c); Wei et al. (2024b); Zhang et al. (2024i). As these multi-encoder models project different modalities into a shared latent space, multimodal RAGs rely on efficient search strategies to retrieve relevant external knowledge.
@@ 4.1.1 Efficient Search and Similarity Retrieval
Maximum inner product search (MIPS) variants are widely used for fast and direct similarity comparisons Tiwari et al. (2024); Wang et al. (2024c); Zhao et al. (2023b). Systems such as MuRAG Chen et al. (2022a) and RA-CM3 Yasunaga et al. (2023) employ approximate MIPS to efficiently retrieve the top-k candidates by maximizing the inner product between the query vector and a large collection of image–text embeddings. Large-scale implementations leverage distributed MIPS techniques, such as TPU-KNN Chern et al. (2022), for high-speed retrieval. Additionally, ScaNN (Scalable Nearest Neighbors) Guo et al. (2020), MAXSIM score Chan and Ng (2008); Cho et al. (2024), and approximate KNN methods Caffagni et al. (2024) have been adopted for efficient similarity computation.
@@
Recent advancements in MIPS optimization focus on reducing retrieval latency and improving accuracy. These include adaptive quantization Zhang et al. (2023a); Li et al. (2024a), hybrid sparse-dense representations Nguyen et al. (2024); Zhang et al. (2024a), and learned index structures Zhai et al. (2023); Basnet et al. (2024). These techniques help optimize retrieval performance by balancing efficiency and precision in multimodal RAG systems.
4.1.2 Modality-Based Retrieval
Modality-aware retrieval techniques optimize retrieval efficiency by leveraging the unique characteristics of each modality.
Text-Centric Retrieval
This remains foundational in multimodal RAG systems, with both traditional methods like BM25 Robertson and Zaragoza (2009), and dense retrievers such as MiniLM Wang et al. (2020a) and BGE-M3 Chen et al. (2024b) dominate text-based evidence retrieval Chen et al. (2022b); Suri et al. (2024); Nan et al. (2024). Novel approaches also address the need for fine-grained semantic matching and domain specificity: For instance, ColBERT Khattab and Zaharia (2020) and PreFLMR Lin et al. (2024b) employ token-level interaction mechanisms that preserve nuanced textual details to improve precision for multimodal queries, while RAFT Zhang et al. (2024h) and CRAG Yan et al. (2024) enhance retrieval performance by ensuring precise citation of text spans.
Vision-Centric Retrieval
This focuses on directly leveraging image representations for knowledge extraction Kumar and Marttinen (2024); Yuan et al. (2023). Systems such as EchoSight Yan and Xie (2024) and ImgRet Shohan et al. (2024) retrieve visually similar content by using reference images as queries. In addition, composed Image Retrieval (CMI) models Feng et al. (2023); Zhao et al. (2024); Jang et al. (2024) enhance retrieval by integrating multiple image features into a unified query representation. Similarly, Pic2word Saito et al. (2023) maps visual content to textual descriptions, enabling zero-shot image retrieval.
Video-Centric Retrieval
@@ 4.1.2 Modality-Based Retrieval
These methods extend vision-based techniques by incorporating temporal dynamics and large video-language models (LVLMs), driven by novel frameworks like iRAG Arefeen et al. (2024), which introduces incremental retrieval for sequential video understanding, and MV-Adapter Jin et al. (2024), optimizing multimodal transfer learning for video-text retrieval. Recent breakthroughs focus on long-context processing: Video-RAG Luo et al. (2024b) leverages visually aligned auxiliary texts (OCR/ASR) to enhance retrieval without proprietary models, T-Mass Wang et al. (2024b) models text as a stochastic embedding to enhance text-video retrieval, and VideoRAG Ren et al. (2025) -employs dual-channel architectures with graph-based knowledge grounding for extreme-length videos. For temporal reasoning, CTCH Shen et al. (2024) uses contrastive transformer hashing to model long-term dependencies, while RTime Du et al. (2024) introduces reversed-video hard negatives to rigorously benchmark temporal causality. Meanwhile, for complex video understanding, OmAgent Zhang et al. (2024e) adopts a divide-and-conquer framework, and DRVideo Ma et al. (2024d) addresses long-video understanding with a document-based retrieval approach.
@@
Recent research has advanced beyond traditional uni-modal retrieval methods toward models that directly process entire document pages. These approaches integrate textual, visual, and structural elements like tables, font styles, and page layouts to enhance retrieval performance in complex documents. ColPali Faysse et al. (2024) pioneers end-to-end document image retrieval by embedding page patches with a vision-language backbone, bypassing OCR entirely. Models like ColQwen2 Wang et al. (2024d); Khattab and Zaharia (2020) and M3DocVQA Cho et al. (2024) extend this paradigm with dynamic resolution handling and holistic multi-page reasoning.
Newer frameworks refine efficiency and layout understanding: ViTLP Mao et al. (2024) and DocLLM Wang et al. (2024a) pre-train generative models to align spatial layouts with text, while CREAM Zhang et al. (2024b) employs coarse-to-fine retrieval with multimodal efficient tuning to balance accuracy and computation costs. Finally, mPLUG-DocOwl 1.5 Hu et al. (2024a) and 2 Hu et al. (2024b) unify structure learning across formats (e.g., invoices, forms) without OCR dependencies.
4.1.3 Re-ranking and Selection Strategies
Effective retrieval in multimodal RAG systems requires not only identifying relevant information but also prioritizing retrieved candidates. Re-ranking and selection strategies enhance retrieval quality by optimizing example selection, refining relevance scoring, and applying filtering mechanisms.
Optimized Example Selection
These techniques often employ multi-step retrieval, integrating both supervised and unsupervised selection strategies Luo et al. (2024a); Yuan et al. (2023). For instance, Su et al. (2024a) enhance multimodal inputs using probabilistic control keywords to improve credibility, RULE Xia et al. (2024b) calibrates retrieved context selection via statistical methods like the Bonferroni correction to mitigate factuality risks, and clustering-based key-frame selection ensures diversity in video-based retrieval Dong et al. (2024b).
Relevance Score Evaluation
Several methods employ advanced scoring mechanisms to improve retrieval relevance Mortaheb et al. (2025b, a); Zhi Lim et al. (2024). Multimodal similarity measures, including structural similarity index measure (SSIM) Wang et al. (2020b), normalized cross-correlation (NCC), and BERTScore Zhang et al. (2020), aid in re-ranking retrieved documents. Cross-encoders trained on sequence classification tasks refine relevance scoring within retrieval pipelines Zhi Lim et al. (2024), while hierarchical post-processing techniques integrate retrieval, passage-level, and answer confidence scores for improved ranking Zhang et al. (2024g); Yan and Xie (2024); Xu et al. (2024a). LDRE Yang et al. (2024) employs semantic ensemble methods to adaptively weigh multiple caption features, whereas RAGTrans Cheng et al. (2024) and OMG-QA Nan et al. (2024) incorporate traditional ranking functions like BM25 Robertson and Zaragoza (2009).
Filtering Mechasnism
This ensures high-quality retrieval by eliminating irrelevant data. Hard negative mining, as used in GME Zhang et al. (2024i) and MM-Embed Lin et al. (2024a), mitigates modality bias through modality-aware sampling and synthesized negatives. Similarly, consensus-based filtering approaches, seen in MuRAR Zhu et al. (2025) and ColPali Faysse et al. (2024), employs source attribution and multi-vector mapping to filter out low-similarity candidates. Moreover, dynamic modality filtering methods, such as RAFT Zhang et al. (2024h), Img2Loc Zhou et al. (2024e), and MAIN-RAG Chang et al. (2024), train retrievers to disregard confusing data, thereby enhancing the discriminative capacity and overall robustness of multi-modal retrieval systems.
@@ 4.2 Fusion Mechanisms
4.2.1 Score Fusion and Alignment
Models in this category utilize distinct strategies to align multimodal representations. Zhi Lim et al. (2024) convert text, tables, and images into a single textual format using a cross-encoder trained for relevance scoring. Sharifymoghaddam et al. (2024) introduce interleaved image–text pairs that vertically merge multiple few-shot images (as in LLaVA Liu et al. (2023a)), while aligning modalities via CLIP score fusion Hessel et al. (2021) and BLIP feature fusion Li et al. (2022a). Riedler and Langer (2024), Wiki-LLaVA Caffagni et al. (2024), C3Net Zhang et al. (2024c), and MegaPairs Zhou et al. (2024a), embed images and queries into a shared CLIP space.
@@ 4.2.1 Score Fusion and Alignment
VISA Ma et al. (2024b) employs the Document Screenshot Embedding (DSE) model to align textual queries with visual document representations by encoding both into a shared embedding space. REVEAL Hu et al. (2023) injects retrieval scores into attention layers to minimize L2-norm differences between query and knowledge embeddings, and MA-LMM He et al. (2024) aligns video-text embeddings via a BLIP-inspired Query Transformer Li et al. (2022a). LLM-RA Jian et al. (2024) concatenates text and visual embeddings into joint queries to reduce retrieval noise, while RA-BLIP Ding et al. (2024b) employs a 3-layer BERT-based adaptive fusion module to unify visual–textual semantics. Xue et al. (2024) use a prototype-based embedding network Zheng et al. (2023) to map object-predicate pairs into a shared semantic space, aligning visual features with textual prototypes. Re-IMAGEN Chen et al. (2022b) balances creativity and entity fidelity in text-to-image synthesis via interleaved classifier-free guidance during diffusion sampling. To improve multimodal alignment, VISRAG Yu et al. (2024) enhances alignment with position-weighted mean pooling on VLM hidden states, prioritizing later tokens for relevance, and RAG-Driver Yuan et al. (2024) aligns visual–language embeddings using visual instruction tuning and an MLP projector.
4.2.2 Attention-Based Mechanisms
Attention-based methods dynamically weight cross-modal interactions to support task-specific reasoning. EMERGE Zhu et al. (2024b), MORE Cui et al. (2024), and AlzheimerRAG Lahiri and Hu (2024) integrate heterogeneous data via cross-attention. RAMM Yuan et al. (2023) employs a dual-stream co-attention transformer, combining self-attention and cross-attention to fuse retrieved biomedical images/texts with input data. RAGTrans Cheng et al. (2024) applies user-aware attention to social media features.